Determine the maximum power transfer to terminals ab 5Ω 10 Ω 25 V 0.51, 10Ω OD
Answer #1

Q Determine maximum power toansfer to terminals ab, 51 Tos 25V (40.55 136 - 06 To solve this type of questions you should fin* write Nodal equation at b. 6-va + 10 10 V6-vatub 10 Qub=va = Va = QV46 th= olio Vb 3 = uth 10 - and ix ohms law 3 substiQua + 10x0.5lx + Va-Vr 10 Bra +5ix =VT - at Node i bl. 6-va 10 IT ub 10 + QV6-va = 10 IT & 10 IT Qur - vai ix VT 10 From equa